OpenStudy (anonymous):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits. Can someone pls explain how to do this?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

@sweetburger

OpenStudy (anonymous):

@vera_ewing

OpenStudy (sweetburger):

well the area of a square is modeled via A=s^2 so your answer would be found from 6.25^2

OpenStudy (sweetburger):

correct significant digits would be 3 digits

OpenStudy (anonymous):

so how do i get the answer?

OpenStudy (sweetburger):

6.25x6.25

OpenStudy (anonymous):

ok and do i have to do anything after that?

OpenStudy (sweetburger):

no just correct sig figs

OpenStudy (anonymous):

ok thnk :)

OpenStudy (anonymous):

i still dont get it lol sorry @chrisdbest

OpenStudy (jacob902):

A≈39.06ft²

OpenStudy (anonymous):

is that the answer?

OpenStudy (jacob902):

yep

OpenStudy (anonymous):

can you pls tell me how you got that answer? :)

OpenStudy (chrisdbest):

Lol, ppl these days

OpenStudy (anonymous):

ok then can you explain it @chrisdbest pls :)

OpenStudy (chrisdbest):

Be like " Dang she cute. She MUST deserve a medal for FIRST PRIZE!" Just Giving Cute girls medals when she don't understand something.

OpenStudy (anonymous):

can you just pls explain thnx

OpenStudy (chrisdbest):

Yeah, a square has 4 equal sides

OpenStudy (chrisdbest):

so each side is 6.25

OpenStudy (chrisdbest):

And Area = L*W

OpenStudy (teddyiswatshecallsme):

All you need to do is plug in S which is 6.25ft So, a = s^2 would turn into a =6.25^2 Then all you do is just punch that into your calculator and whatever you get is your answer. Then that's just a = s

OpenStudy (sweetburger):

literally almost word for word what i said^

OpenStudy (chrisdbest):

So you do 6.25 * 6.25

OpenStudy (chrisdbest):

And BAM! You get 39.0625

OpenStudy (anonymous):

yea i got that lol so thats it?

OpenStudy (teddyiswatshecallsme):

Yes Ma'am.

OpenStudy (anonymous):

thnx guys :)